### v3

# Learning to Bid in Non-Stationary Repeated First-Price Auctions

[…]
Learning to Bid in Non-Stationary Repeated First-Price Auctions
[…]
Stern School of Business, New York University2
Department of IEDA, The Hong Kong University of Science and Technology3
[…]
First-price auctions have recently gained significant traction in digital advertising markets, exemplified by Google’s transition from second-price to first-price auctions. Unlike in second-price auctions, where bidding one’s private valuation is a dominant strategy, determining an optimal bidding strategy in first-price auctions is more complex. From a learning perspective, the learner (a specific bidder) can interact with the environment (other bidders, i.e., opponents) sequentially to infer their behaviors. Existing research often assumes specific environmental conditions and benchmarks performance against the best fixed policy (static benchmark). While this approach ensures strong learning guarantees, the static benchmark can deviate significantly from the optimal strategy in environments with even mild non-stationarity. To address such scenarios, a dynamic benchmark—representing the sum of the highest achievable rewards at each time step—offers a more suitable objective. However, achieving no-regret learning with respect to the dynamic benchmark requires additional constraints. By inspecting reward functions in online first-price auctions, we introduce two metrics to quantify the regularity of the sequence of opponents’ highest bids, which serve as measures of non-stationarity. We provide a minimax-optimal characterization of the dynamic regret for the class of sequences of opponents’ highest bids that satisfy either of these regularity constraints. Our main technical tool is the Optimistic Mirror Descent (OMD) framework with a novel optimism configuration, which is well-suited for achieving minimax-optimal dynamic regret rates in this context. We then use synthetic datasets to validate our theoretical guarantees and demonstrate that our methods outperform existing ones.
[…]
Learning to Bid; Online First-price Auctions; Non-stationary Online Learning.
1 Introduction
By 2029, global digital advertising spending is projected to reach $1126$ billion (Statista 2023). As online ad display grows in importance, it has become a central focus in operations research, information systems, and machine learning (see e.g., Wang et al. 2017, Choi et al. 2020). In online ad markets (also known as ad exchanges), advertisers bid for ad impressions offered by publishers on ad exchanges through auctions to maximize their rewards, while publishers manage inventory to optimize customer impressions. Specifically, in each auction round, the publisher displays an ad impression to potential advertisers (buyers), who assess its value and submit bids. Allocation and pricing of impressions are then determined by an online auction protocol.
Second-price auctions, championed by Nobel-prize-winning work of Vickrey (1961), have been widely used in online ad markets (Edelman et al. 2007, Despotakis et al. 2021) for their incentive compatibility, which encourages truthful bidding. In this format, the highest bidder wins the ad impression but pays the second-highest bid. Despite its theoretical elegance, second-price auctions face practical criticisms, particularly the potential for auctioneers to manipulate the second-highest bid to inflate payments undetectably (Rothkopf et al. 1990, Lucking-Reiley 2000, Akbarpour and Li 2020). In online ad auctions, such manipulation allows ad exchanges to substantially increase revenue. Concerns over trust and the rise of publisher-initiated header bidding (Despotakis et al. 2021) have led major ad exchanges—including Google AdSense (Wong 2021), Google Ad Manager (Bigler 2019), Yahoo Advertising (Alcobendas and Zeithammer 2021), and Xandr (Microsoft Learn Challenge 2024)—to shift to first-price auctions. In these, the highest bidder wins and pays their bid, thereby addressing trust issues. However, first-price auctions lack incentive compatibility, as revealing a bidder’s true valuation is no longer optimal. This raises a critical question: what bidding strategies should bidders adopt in online first-price auctions to maximize cumulative rewards?
Two primary perspectives address this problem: the game-theoretic and online learning approaches. From the game-theoretic perspective, the problem originates from foundational work by Vickrey (1961), Myerson (1981), where bidders are modeled as rational agents with partial or complete information about competitors’ valuation distributions. This framework allows derivation of optimal strategies and Bayesian Nash equilibria. While significant progress has been made in computing Bayesian Nash equilibria for online first-price auctions (Wang et al. 2020, Filos-Ratsikas et al. 2021, Bichler et al. 2023, Chen and Peng 2023, Filos-Ratsikas et al. 2024), these methods often assume bidders have some precise knowledge of valuation distributions. While such assumptions may hold in physical auctions—where industry peers have insights into each other’s valuations—they are less realistic in online auctions, where bidders typically lack information about competitors’ identities, making valuation estimation far more challenging.
An alternative is the online learning perspective, where a specific bidder is treated as the “learner” and the remaining bidders are modeled as the “environment” (potentially with some assumptions on the environment to ensure learnability). In this view, the problem of finding the optimal bidding strategy can be cast as a sequential two-player game. At the start of the game, the learner is assumed to have no knowledge of the environment. However, based on past decisions and the feedback received, the learner can iteratively update their bidding strategy. A common performance metric in this perspective is called (static) regret, which measures the difference between the cumulative reward achieved by the best fixed policy and the cumulative reward of the learning algorithm. The goal of online learning is to achieve sublinear regret, which ensures that the time-averaged performance of the learning algorithm asymptotically converges to that of the best fixed policy. This perspective has inspired a body of seminal work (Han et al. 2020, Zhang et al. 2022, Badanidiyuru et al. 2023, Balseiro et al. 2023, Han et al. 2025) focused on achieving sublinear regret in both stochastic and adversarial settings. In these contexts, the private value of the bidder and/or the opponents’ highest bid at each time step are either independently and identically distributed (i.i.d.) or adversarially generated.
While these approaches provide strong theoretical guarantees, real-world scenarios are usually much more involved. The value of a fixed ad impression can change over time, possibly exhibiting seasonal periodic trends or even sudden shifts due to unforeseen events. The values of different ad impressions can be correlated in complex ways, since advertisers may have complementary marketing campaigns, competing objectives, or overlapping target audiences. For a certain learner, their opponents’ bidding strategies may also evolve over time, as they adapt to the learner’s bidding behavior. These complexities often fall outside the assumptions of purely stochastic or adversarial environments, and it is more natural to term these situations as non-stationary environments. In such environments, considering competing with the best fixed policy (static benchmark) might oversimplify the situation that the learner faces. In contrast, a dynamic benchmark—representing the maximum achievable cumulative reward—is always optimal, even in non-stationary settings. Learning in non-stationary environments poses a fundamental challenge in operations research (Besbes et al. 2015, 2019, Cheung et al. 2022, 2023) and machine learning (Yang et al. 2016, Zhang et al. 2018, Wei and Luo 2021), yet it remains underexplored in the context of online first-price auctions. This gap motivates the core focus of this work, which investigates the following key questions:
For online first-price auctions in non-stationary ad markets, can we effectively compete with a dynamic benchmark? What mathematical tools can help establish minimax-optimal dynamic regret rates in such settings?
1.1 Our Contributions
Consider a learner participating in a $T$-round online first-price auction. In each round $t$, the learner observes an ad impression, receives a private valuation $v_{t}$, and then determines a bid $b_{t}$. After submitting $b_{t}$, the learner observes $m_{t}$, the highest bid from other participants, and receives a reward of $(v_{t}-b_{t})\cdot\mathbbm{1}(b_{t}\geq m_{t})$. This full-information feedback setting is widely used in practice, including Google Ad Exchange (Google Developers 2024). Other feasible feedback types include binary feedback (Balseiro et al. 2023) and winning-bid feedback (Han et al. 2025). Our main contributions are summarized as follows:
  • Inspired by previous work on non-stationary online learning (Besbes et al. 2015, Jadbabaie et al. 2015, Yang et al. 2016), we propose two regularity conditions on the sequence of opponents’ highest bids to characterize the extent of non-stationarity: the temporal variation $V_{T}\coloneqq\sum_{t=2}^{T}|m_{t}-m_{t-1}|$ and the number of abrupt switches $L_{T}\coloneqq\sum_{t=2}^{T}\mathbbm{1}(m_{t}\neq m_{t-1})$. If either $V_{T}=\Omega(T)$ or $L_{T}=\Omega(T)$, then any non-anticipatory policy suffers $\Omega(T)$ dynamic regret. Thus, a reasonable goal is to achieve sublinear dynamic regret rates when $V_{T}=o(T)$ or $L_{T}=o(T)$. Notably, these regularity conditions do not depend on the learner’s private valuation sequence, allowing the learner’s private valuation sequence to be adversarially generated.
[…]
  • We establish $\Omega(\sqrt{TV_{T}})$ and $\Omega(\sqrt{L_{T}})$ minimax lower bounds for online first-price auction instances regularized by either $V_{T}$ or $L_{T}$, respectively. For sequential learning of convex and Lipschitz functions with exact feedback, the dynamic regret lower bound in terms of $V_{T}$ is $\Omega(V_{T})$ (Jadbabaie et al. 2015, Yang et al. 2016). Our results, therefore, highlight a sharp separation between learning one-sided Lipschitz functions and convex, Lipschitz functions. To prove the lower bounds, we construct batches with small temporal variations. Within each batch, the optimal dynamic regret of any non-anticipatory policy can be computed via dynamic programming. By suitably concatenating these batches, we derive the desired lower bounds.
  • Since both $V_{T}$ and $L_{T}$ capture different types of regularity in the opponents’ highest bid sequence, it is desirable to achieve a best-of-both-worlds guarantee of $\tilde{O}(\min\{\sqrt{TV_{T}},L_{T}\})$, which automatically adapts to the better of the two bounds. We achieve this by combining our algorithms using the meta algorithm of Sani et al. (2014). The theoretical guarantees are summarized in Table 1. Notably, the lower bounds hold even if $V_{T}$ or $L_{T}$ is known in advance, while the upper bounds do not require such prior knowledge.
[…]
    Regularity Upper Bound Lower Bound
    $V_{T}$ $\tilde{O}(\sqrt{TV_{T}})$, Theorem 4.6 $\Omega(\sqrt{TV_{T}})$, Theorem 5.2
    $L_{T}$ $\tilde{O}(L_{T})$, Theorem 4.9 $\Omega(L_{T})$, Theorem 5.4
    Best-of-both-worlds $\tilde{O}(\min\{\sqrt{TV_{T}},L_{T}\})$, Theorem 4.10 -
  • We evaluate our theoretical findings using synthetic datasets. We first confirm that our algorithms achieve the theoretical dynamic regret rates. We then consider a multi-agent bidding environments where the opponents of the learner run the budget-pacing policy by Gaitonde et al. (2022), and demonstrate that our algorithms outperform two important baselines: the Hedge algorithm and the SEW policy (Han et al. 2020), especially in the regime where opponents have limited budgets.
1.2 Key Challenges
The primary challenge in proving the upper bounds stems from the one-sided Lipschitz property of the reward function in online first-price auctions. In simple terms, bidding slightly higher than necessary results in only a minor revenue loss, while bidding slightly lower can cause a much larger loss. To address non-stationarity in the data, we adopt the restart scheme introduced by Besbes et al. (2015)—dividing the time horizon $T$ into batches and restarting a dedicated algorithm at the beginning of each batch. Our analysis decomposes the dynamic regret into two components: the static regret and the transition cost that bridges static and dynamic regret. Unlike previous work, we address these two terms using novel analytical tools. For instance, to bound the transition cost as in Besbes et al. (2015), one needs to bound the temporal variation of the reward sequence by that of the opponents’ highest bid sequence, which roughly amounts to bounding the variation of rewards by the variation of the maximizers [^1] The correspondence is not exact: when $v_{t}\geq m_{t}$, the opponents’ highest bid $m_{t}$ is the reward maximizer, but when $v_{t}<m_{t}$, the reward is maximized for any bid smaller than $m_{t}$.. While previous literature (Jadbabaie et al. 2015, Yang et al. 2016) assumes full Lipschitzness, we relax this assumption by relying solely on the one-sided Lipschitz property.
To control the static regret, we recall that the SEW policy (Han et al. 2020) achieves the minimax-optimal $\tilde{O}(\sqrt{T})$ static regret bound (Han et al. 2025) in online first-price auctions with adversarial inputs. Moreover, in settings with convex Lipschitz loss functions and noisy feedback, Besbes et al. (2015) demonstrated that restarting the online gradient descent (OGD) algorithm, which achieves minimax-optimal static regret, ensures minimax-optimal dynamic regret. Surprisingly, directly plugging the SEW policy into the restart scheme does not produce the same optimal dynamic regret rates— likely because the SEW policy, designed for adversarial environments, lacks adaptivity in slowly varying settings.
To improve the sublinear static regret guarantee under slow variation, we employ Optimistic Mirror Descent (OMD) (Chiang et al. 2012, Rakhlin and Sridharan 2013, Wei and Luo 2018)—a variant of mirror descent that incorporates an “optimistic” guess of the gradient of the expected reward for the current round. When this guess is taken as the gradient from the previous round, OMD can yield much lower static regret than $O(\sqrt{T})$ in slowly varying environments. Yet, no previous work has used OMD to reach minimax-optimal dynamic regret rates. By carefully configuring the optimism vector, we show that OMD’s static regret can be bounded by the transition cost plus a small additive term, effectively balancing the trade-off between static regret and transition cost and leading to minimax-optimal dynamic regret rates.
The difficulty in proving the lower bounds arises because the learner directly observes the highest bids of the others rather than receiving noisy feedback. Noisy feedback simplifies many information-theoretic arguments—such as those based on Le Cam’s method—which have been crucial for deriving lower bounds in non-stationary online learning (Besbes et al. 2015, 2019, Cheung et al. 2022) and in online first-price auctions (Han et al. 2020, Zhang et al. 2022, Cesa-Bianchi et al. 2024). In our setting, alternative approaches are required. Our optimal lower bounds are inspired by the minimax lower bounds for learning with a few experts (Cover 1966, Gravin et al. 2016, Harvey et al. 2023), which establish lower bounds by constructing suitable problem instances whose minimax value can be evaluated. Specifically, leveraging the one-sided Lipschitz property, we construct batches of opponents’ highest bids with small temporal variation, but any non-anticipatory policy suffers a large amount of dynamic regret in these batches. We then carefully stitch these batches together to obtain the desired lower bound results.
1.3 Paper Organization
The paper is organized as follows. In Section 2, we review prior work, positioning our contributions within the existing literature. Section 3 formally defines the problem setting and introduces our methodology. In Section 4, we present our upper bound results by deriving dynamic regret rates within the Optimistic Mirror Descent (OMD) framework. Section 5 then provides lower bounds via a minimax analysis, thereby establishing the optimality of our upper bounds. Section 6 offers numerical simulations that validate the dynamic regret rates of our proposed algorithms and compare them with baseline approaches in a multi-agent bidding environment. Finally, Section 7 summarizes our findings and discusses potential directions for future research.
1.4 Notations
Let $v_{t}$ and $m_{t}$ denote the learner’s private valuation and the highest bid from other bidders at round $t$, respectively. We denote the learner’s bid at round $t$ by $b_{t}$. Following previous work (Han et al. 2020, Balseiro et al. 2023, Han et al. 2025), we assume $v_{t},m_{t},b_{t}\in[0,1]$. Since the other bidders’ highest bid $m_{t}$ is observed by the learner, we discretize the decision space $[0,1]$ into $N$ discrete bidding prices and model the problem as learning with expert advice (Cesa-Bianchi and Lugosi 2006) with $N$ experts. Let $r_{t,i}$ denote the reward of the $i$-th expert at round $t$.
[…]
2 Related Work
In this section, we briefly review relevant work on first-price auctions and online learning in non-stationary environments.
First-price Auctions.
Although Vickrey is more commonly associated with the second-price auction, Vickrey (1961) formalize and compare several auction formats, including the first-price auction. In recent years, as certain online ad exchanges switch from second-price to first-price auctions, first-price auctions gain increasing attention from researchers in economics, operations research, and machine learning. From a game-theoretic perspective, researchers study aspects such as the Bayesian Nash equilibrium, pacing equilibrium, and algorithmic collusion behaviors in first-price auctions (Wang et al. 2020, Filos-Ratsikas et al. 2021, Conitzer et al. 2022, Banchio and Skrzypacz 2022, Banchio and Mantegazza 2023, Chen and Peng 2023, Bichler et al. 2023, Jin and Lu 2023, Balseiro et al. 2023).
This work focuses on a learning perspective, where a learner sequentially interacts with the environment to learn an optimal bidding strategy. Inspired by patterns in real-world auction data, Zhang et al. (2021) introduce a non-parametric approach for bid updates, demonstrating its superiority over traditional parametric methods. Balseiro et al. (2023) employ cross-learning to improve regret rates for online first-price auctions with binary feedback. When $v_{t}$ is i.i.d. from a known distribution and $m_{t}$ is chosen adversarially, they achieve a regret rate of $\tilde{O}(T^{\frac{2}{3}})$, improving upon the $\tilde{O}(T^{\frac{3}{4}})$ rate achieved by standard contextual bandit techniques. Later, Schneider and Zimmert (2024) extend these results to the setting where the distribution of $v_{t}$ is unknown, achieving the same $\tilde{O}(T^{\frac{2}{3}})$ regret through novel techniques. Han et al. (2020) study online first-price auctions with full-information feedback when both $v_{t}$ and $m_{t}$ are chosen adversarially. Using the tree-chaining technique (Cesa-Bianchi et al. 2017), they achieve a regret rate of $\tilde{O}(\sqrt{T})$ against the set of 1-Lipschitz policies. When $m_{t}$’s are i.i.d. generated, Han et al. (2025) improve the analysis of Balseiro et al. (2023) to the winning-bid feedback setting throught some novel observations, demonstrating $\tilde{O}(\sqrt{T})$ regret. Additionally, Zhang et al. (2022) explore improved regret guarantees by incorporating hints about bidding profiles. Badanidiyuru et al. (2023) consider online first-price auctions where $m_{t}$ is generated by a context vector with log-concave noise and establish $\tilde{O}(\sqrt{T})$ regret guarantees under full-information feedback. Wang et al. (2023) investigate first-price auctions with budget constraints, achieving sublinear regret rates when both $v_{t}$ and $m_{t}$ are i.i.d. Kumar et al. (2024) study settings where $v_{t}$ is i.i.d. and $m_{t}$ is adversarially chosen, achieving $\tilde{O}(\sqrt{T})$ regret that is both rate-optimal and strategically robust. Cesa-Bianchi et al. (2024) characterize minimax-optimal static regret rates for various feedback settings, highlighting the role of auction format transparency. All the aforementioned works focus on competing against the best fixed policy within a pre-determined policy set, whereas our work aims to compete with the policy that achieves the maximum possible revenue.
Recently, there is a growing body of work on online first-price auctions, where the learner faces additional constraints such as budget or ROI constraints. Balseiro and Gur (2019) propose the budget-pacing dynamics in online second-price auctions, which use a sequence of Lagrangian multipliers to shade the learner’s bid. Gaitonde et al. (2022) generalize this idea to online first-price auctions with budgets, while Lucier et al. (2024) further allow the existence of ROI constraints. Other related work includes Ai et al. (2022), Castiglioni et al. (2022), Wang et al. (2023), Fikioris and Tardos (2023), Aggarwal et al. (2025). Both Gaitonde et al. (2022) and Lucier et al. (2024) consider to compete with a dynamic benchmark as well, but there are some key differences between their work and ours. First, Gaitonde et al. (2022), Lucier et al. (2024) consider value maximizing bidders while our work considers revenue maximizing bidders; the value maximizing bidders make sense in the constrained setting but not for the unconstrained setting, since the bidder has the incentive to win every ad impression. Second, Gaitonde et al. (2022), Lucier et al. (2024) consider competing with a sequence of Lagrangian multipliers, where each Lagrangian multiplier makes the expected expenditure at that round equal to the ratio of the initial budget and the time horizon. This sequence is not guaranteed to achieve the highest possible cumulative value, so the dynamic benchmark considered therein is weaker than ours. Third, it is unknown whether Gaitonde et al. (2022), Lucier et al. (2024) achieve minimax-optimal regret rates even with respect to this relaxed dynamic benchmark notion.
Learning in Non-stationary Environments.
Besbes et al. (2015) study stochastic optimization in non-stationary environments, where the loss at each round may vary, and show that sublinear dynamic regret is achievable when the temporal variation—a measure of the total change in the loss function over time—is sublinear in the time horizon. Besbes et al. (2015) provide minimax-optimal characterizations of dynamic regret for online convex optimization and bandit convex optimization. Our problem formulation is inspired by Besbes et al. (2015), but we make necessary adjustments to better accommodate the one-sided Lipschitzness of the reward function. Please see Remark 3.2 in Section 3.1 for a comprehensive comparison. Besbes et al. (2015) assumes that the temporal variation of the loss sequence is known in advance. Jadbabaie et al. (2015) demonstrate how to remove this assumption in the online convex optimization setting. Additionally, Jadbabaie et al. (2015), Yang et al. (2016), Zhang et al. (2018), Baby and Wang (2021, 2022) explore alternative definitions of dynamic regret, such as the path-length of the minimizers of the loss functions, and establish corresponding dynamic regret guarantees. These formulations are also related to ours, but these approaches rely on strong convexity, exponential concavity, convexity, Lipschitzness or smoothness of the loss functions, while the reward function we consider is merely one-sided Lipschitz. Besbes et al. (2019) investigate multi-armed bandit problems under non-stationary reward distributions, demonstrating that sublinear regret can be achieved if the total variation of these distributions is known and sublinear in the time horizon. To remove the need for prior knowledge of the variation budget, Cheung et al. (2022) propose the bandit-over-bandit technique, which applies to various non-stationary stochastic bandit problems. Building on this, Zhao et al. (2021) simplify the analysis in Cheung et al. (2022) and derive sublinear regret bounds for linear bandits with variable decision sets. In the context of reinforcement learning (RL), Cheung et al. (2023) employ a similar bandit-over-RL approach to tackle non-stationary settings, achieving nearly optimal regret bounds. Wei and Luo (2021) provide a general framework for non-stationary online learning, covering both linear bandits and RL, and achieve optimal dynamic regret rates. Simchi-Levi et al. (2023) study experimental design under non-stationary linear trends, while Chen et al. (2025) focus on non-stationary multi-armed bandits with periodic mean rewards. Huang and Wang (2025) consider non-stationary online learning with noisy realization of the losses, and achieve minimax-optimal regret guarantees when losses are strongly convex or merely Lipschitz. Though Zhao and Chen (2020) study online second-price auctions in non-stationary settings, their objective and methods differ significantly from ours.
3 Problem Formulation and Main Results
In this section, we introduce the problem formulation for online first-price auctions in non-stationary environments, outline the main algorithmic framework we will use, present the informal main results, and define the notations that will be used throughout the paper.
3.1 Problem Formulation
In non-stationary environments, an advertiser’s valuation for an ad impression can vary over time, requiring advertisers to account for this variability when participating in online first-price auctions. We begin with a general description of the online first-price auction (Han et al. 2020, 2025, Cesa-Bianchi et al. 2024), followed by a formal definition of the dynamic benchmark and possible ways to quantify the degree of non-stationarity.
In this auction format, a set of bidders (advertisers) competes to purchase ad impressions from a publisher. Each round, the publisher displays an ad impression along with relevant details, such as user demographics, keywords, and the ad’s size and location. Each bidder estimates the value of the ad impression and submits a bid. Under the first-price auction protocol, the bidder who offers the highest bid wins the ad impression and pays the bid amount. Formally, the online first-price auction is a game spanning $T$ rounds. In each round $t=1,\dots,T$, the bidder observes an ad impression, generates a private value $v_{t}\in[0,1]$, and submits a bid $b_{t}\in[0,1]$. Let $m_{t}\in[0,1]$ represent the highest bid among other bidders. The bidder’s payoff is then given by
$r(b_{t};v_{t},m_{t})\coloneqq(v_{t}-b_{t})\cdot\mathbbm{1}(b_{t}\geq m_{t}).$
Here, $\mathbbm{1}(b_{t}\geq m_{t})$ is the indicator function that equals $1$ if the bidder wins the auction (i.e., $b_{t}\geq m_{t}$), and 0 otherwise. The one-sided Lipschitz property means: when moving from a higher bid $b^{\prime}\leq v_{t}$ to a slightly lower bid $b$, the reward increase is bounded by the bid difference, but the reward decrease can be significant (particularly when crossing the discontinuity at $m_{t}$).
For simplicity, we assume the time horizon $T$ is known to the learner. If $T$ is unknown, the doubling trick (Auer et al. 2002, Cesa-Bianchi and Lugosi 2006) can be used to eliminate this requirement. Since this is a sequential decision-making problem, it is essential to formally define the information received by the learner before submitting $b_{t}$. We mainly consider the case where the learner observes $m_{t}$, the highest bid from other bidders, so the information up to time t1t-1 can be described by the following filtration:
$\mathcal{H}_{t}\coloneqq\sigma((v_{s},m_{s})_{s=1}^{t-1},v_{t}),$
where $\sigma(\cdot)$ is the $\sigma$-algebra generated by the observations. Conventionally, the filtration up to t1t-1 should not include $v_{t}$, as this represents information from the current round. However, prior work (Han et al. 2020, Balseiro et al. 2023, Han et al. 2025) assumes that the bidder knows $v_{t}$ before determining their bid $b_{t}$. This assumption is reasonable because ad exchanges typically display the ad impression and related contextual or demographic information to bidders, enabling them to estimate the value of the impression. Therefore, we include $v_{t}$ in the filtration.
Let $(\Omega,\mathcal{F},\mathbb{P})$ be a probability space with sample space $\Omega=[0,1]$ and $\sigma$-algebra $\mathcal{F}$. Let $U$ be a random variable defined on this probability space, i.e., $U:\Omega\to\mathbb{R}$ is an $\mathcal{F}$-measurable function. We define the set of admissible policies $\Pi$ as follows:
For each $t\in[T]={1,2,\dots,T}$, let $\pi_{t}:\mathbb{R}^{2t-1}\times\Omega\to\mathbb{R}$ be a measurable function. A policy $\pi\in\Pi$ is then a sequence of such measurable functions: $\pi=(\pi_{1},\pi_{2},\dots,\pi_{T})$. Given a policy $\pi\in\Pi$, the bid $b_{t}$ at time $t$ is determined by:
$b_{t}=\pi_{t}((v_{s},m_{s})_{s=1}^{t-1},v_{t},U).$
Thus, the set of admissible policies $\Pi$ is characterized by the collection of these measurable functions $\{\pi_{t}\}_{t=1}^{T}$. Note that the probability measure $\mathbb{P}$ plays a crucial role in determining the distribution of the random variable $U$ and consequently influences the stochasticity of the bidding process.
Previous work on online first-price auctions typically aims to achieve sublinear regret over $T$ rounds against the best fixed policy in hindsight. Formally, this involves designing a policy to minimize the regret:
$\mathbb{E}[\text{R}_{T}(\pi)]\coloneqq\sup_{f\in\tilde{\mathcal{F}}}\sum_{t=1}^{T}r(f(v_{t});v_{t},m_{t})-\sum_{t=1}^{T}\mathbb{E}[r(b_{t};v_{t},m_{t})],$
where $\tilde{\mathcal{F}}$ is a class of policies. Common choices for $\tilde{\mathcal{F}}$ include the set of $1$-Lipschitz policies (Han et al. 2020), the set of monotone policies (Han et al. 2020) or the set of policies that map $C$ possible valuations to $K$ discrete bids (Balseiro et al. 2023, Schneider and Zimmert 2024).
Here, we refer to $\sup_{f\in\tilde{\mathcal{F}}}\sum_{t=1}^{T}r(f(v_{t});v_{t},m_{t})$ as the static benchmark. In contrast, we define the dynamic benchmark as:
$\sum_{t=1}^{T}r(b_{t}^{*};v_{t},m_{t})\coloneqq\sum_{t=1}^{T}\max_{b\in[0,1]}r(b;v_{t},m_{t})=\sum_{t=1}^{T}\max\{v_{t}-m_{t},0\},$ (1)
[…]
$b_{t}^{*}=\begin{cases}m_{t},&v_{t}\geq m_{t}\\ v_{t},&v_{t}<m_{t}.\end{cases}$ (2)
This sequence achieves optimal revenue while eliminating the ambiguity of the optimal bid.
It is immediate to see that the dynamic benchmark represents the maximum possible revenue that the learner can achieve. Moreover, the dynamic benchmark can outperform the static benchmark by $\Omega(T)$, even in instances of online first-price auctions with mild regularity in the opponents’ highest bid sequence. Example 3.1 illustrates the reason for this discrepancy between static and dynamic benchmarks.
Example 3.1
Assume $v_{t}\equiv 1$ for $t\in[T]$ and
$m_{t}=\begin{cases}0,&1\leq t\leq\frac{T}{2},\\ \frac{1}{2},&\frac{T}{2}+1\leq t\leq T.\end{cases}$
Then
$\begin{split}&\sum_{t=1}^{T}r(b_{t}^{*};v_{t},m_{t})-\sup_{f\in\tilde{\mathcal{F}}}\sum_{t=1}^{T}r(f(v_{t});v_{t},m_{t})\\ =&\sum_{t=1}^{T}\max\{v_{t}-m_{t},0\}-\sup_{f\in\tilde{\mathcal{F}}}\sum_{t=1}^{T}r(f(v_{t});v_{t},m_{t})=\frac{3T}{4}-\frac{T}{2}=\frac{T}{4}.\end{split}$
The main fact we rely on is that $f(v_{t})\equiv f(1)$ can only take a single real value and, as such, cannot be optimal on both segments.
Consequently, a no-regret online learning policy, while converging to the best fixed policy in the long run, does not converge to the policy with the highest possible revenue. In this work, we consider minimizing the following dynamic regret in online first-price auctions:
$\mathbb{E}[\text{DR}_{T}(\pi)]\coloneqq\sum_{t=1}^{T}r(b_{t}^{*};v_{t},m_{t})-\sum_{t=1}^{T}\mathbb{E}[r(b_{t};v_{t},m_{t})].$ (3)
It is well-established (e.g., Besbes et al. (2015), Yang et al. (2016), Zhang et al. (2018), Besbes et al. (2019)) that achieving sublinear dynamic regret uniformly is impossible without imposing further constraints on the problem instances. To ensure no-regret online learning, we investigate policies with sublinear dynamic regret guarantees under the assumption that the regularity of the opponents’ highest bid sequence is sublinear in the time horizon $T$. We consider two specific metrics to quantify this regularity:
$V_{T}\coloneqq\sum_{t=2}^{T}|m_{t}-m_{t-1}|$ (4)
$L_{T}\coloneqq\sum_{t=2}^{T}\mathbbm{1}(m_{t}\neq m_{t-1}),$ (5)
where $V_{T}$ measures the temporal variation of the opponents’ highest bid sequence, while $L_{T}$ measures the number of abrupt switches in the opponents’ highest bid sequence.
Remark 3.2
The regularity conditions on the opponents’ highest bid sequence (Equations (4) and (5)) are inspired by Besbes et al. (2015), where the authors use the temporal variation of reward/loss functions as a regularity measure. In our setting, their measure translates to $\sum_{t=2}^{T}\sup_{b\in[0,1]}|r(b;v_{t},m_{t})-r(b;v_{t-1},m_{t-1})|$. However, we argue that $\sum_{t=2}^{T}|m_{t}-m_{t-1}|$ is a more compact and reasonable metric for measuring non-stationarity. By Proposition 3.3, $\sum_{t=2}^{T}|m_{t}-m_{t-1}|$ is at most twice $\sum_{t=2}^{T}\sup_{b\in[0,1]}|r(b;v_{t},m_{t})-r(b;v_{t-1},m_{t-1})|$. In general, however, $\sum_{t=2}^{T}\sup_{b\in[0,1]}|r(b;v_{t},m_{t})-r(b;v_{t-1},m_{t-1})|$ can be much larger than $\sum_{t=2}^{T}|m_{t}-m_{t-1}|$, as demonstrated in Examples 3.4 and 3.5. The disadvantages of $\sum_{t=2}^{T}\sup_{b\in[0,1]}|r(b;v_{t},m_{t})-r(b;v_{t-1},m_{t-1})|$ stem from: (i) this metric neglects the one-sided Lipschitzness of the reward function; (ii) this metric depends on the sequence $(v_{t})_{t=1}^{T}$, which is unnecessary upon careful inspection.
In contrast, our measure defined in Equation (4) compactly captures the regularity of the opponents’ highest bid sequence while avoiding both disadvantages. Additionally, Besbes et al. (2015, Figure 1) emphasize two types of temporal patterns: continuous change and discrete shocks, which directly correspond to our regularity conditions in Equations (4) and (5), respectively.
Proposition 3.3
For any $v_{t-1},v_{t},m_{t-1},m_{t}\in[0,1]$,
[…]
Example 3.4
Assume $v_{t}\equiv 1$ for $t\in[T]$, and
$m_{t}=\begin{cases}0,\quad t\text{ is odd}\\ \epsilon,\quad t\text{ is even}.\end{cases}$
Then $\sum_{t=2}^{T}|m_{t}-m_{t-1}|=(T-1)\epsilon$ while $\sum_{t=2}^{T}\sup_{b\in[0,1]}|r(b;v_{t},m_{t})-r(b;v_{t-1},m_{t-1})|=T-1$.
Example 3.5
Assume $m_{t}\equiv c$ for $t\in[T]$ where $c\in[0,1]$, and
$v_{t}=\begin{cases}0,\quad t\text{ is odd}\\ 1,\quad t\text{ is even}.\end{cases}$
Then $\sum_{t=2}^{T}|m_{t}-m_{t-1}|=0$ while $\sum_{t=2}^{T}\sup_{b\in[0,1]}|r(b;v_{t},m_{t})-r(b;v_{t-1},m_{t-1})|=T-1$.
We aim to establish bounds on the dynamic regret rates under two different regularity conditions. Formally, we consider the suprema of the expected dynamic regret over two sets of feasible opponents’ highest bid sequences $\sup_{(v_{t},m_{t})_{t=1}^{T}\in\mathcal{V}}\mathbb{E}[\text{DR}_{T}(\pi)]$ and $\sup_{(v_{t},m_{t})_{t=1}^{T}\in\mathcal{L}}\mathbb{E}[\text{DR}_{T}(\pi)]$ where the sets $\mathcal{V}$ and $\mathcal{L}$ are defined as follows:
$\mathcal{V}=\left\{\{(v_{t},m_{t})_{t=1}^{T}\}:\sum_{t=2}^{T}|m_{t}-m_{t-1}|\leq V_{T}\right\},\quad\mathcal{L}=\left\{\{(v_{t},m_{t})_{t=1}^{T}\}:\sum_{t=2}^{T}\mathbbm{1}(m_{t}\neq m_{t-1})\leq L_{T}\right\}.$
Here, $\mathcal{V}$ represents the set of opponents’ highest bid sequences with variation bounded by $V_{T}$, while $\mathcal{L}$ represents the set of opponents’ highest bid sequences with a limited number of changes, bounded by $L_{T}$. Before establishing dynamic regret rates, we first present a result that highlights the necessity of assuming sublinear regularity in the time horizon.
Proposition 3.6
Assume $c_{1}\in\left[0,\frac{1}{2}\right]$, then
  • $V_{T}\geq c_{1}T$ implies
    $\inf_{\pi\in\Pi}\sup_{(v_{t},m_{t})_{t=1}^{T}\in\mathcal{V}}\mathbb{E}\left[\textnormal{DR}_{T}(\pi)\right]\geq c_{1}T$
    holds for any admissible policy.
  • $L_{T}\geq c_{1}T$ implies
    $\inf_{\pi\in\Pi}\sup_{(v_{t},m_{t})_{t=1}^{T}\in\mathcal{L}}\mathbb{E}\left[\textnormal{DR}_{T}(\pi)\right]\geq c_{1}^{2}T$
    holds for any admissible policy.
Based on Proposition 3.6, a reasonable objective is to achieve sublinear dynamic regret guarantees when either $V_{T}=o(T)$ or $L_{T}=o(T)$. We establish the corresponding upper bounds and lower bounds in Sections 4 and 5, respectively.
3.2 The Optimistic Mirror Descent Framework
For online first-price auctions, the learner intends to determine a bid $b_{t}\in[0,1]$ for each round $t$. For the convenience of algorithmic implementation, we discretize the interval $[0,1]$ into $N$ discrete candidate bidding prices and maintain a probability distribution $p_{t}$ that governs the probability of selecting the $i$-th discrete bidding price. We can then dynamically adjust the probability of these prices based on their historical performance. Readers familiar with online learning will recognize that we are considering the learning with expert advice framework (Cesa-Bianchi and Lugosi 2006), where each expert suggests a potential bidding price.
A key challenge in non-stationary online learning is that the reward sequence may exhibit continuous drifts or abrupt shifts, so the learner might want to adapt more to the local trend. Our algorithms are composed of two ingredients: a restart scheme, pioneered by Besbes et al. (2015), which decomposes the time horizon into batches satisfying certain criteria; and a static regret minimizer, which is applied to each batch. The static regret minimizer we use can be considered as instantiations of the Optimistic Mirror Descent (OMD) framework developed by Chiang et al. (2012), Rakhlin and Sridharan (2013), Syrgkanis et al. (2015).
The OMD framework (as shown in Algorithm 1) is a two-stage online mirror descent algorithm. At the beginning of each round $t$, the reward vector $r_{t}$ is not yet available to the learner, so the learner adopts the first online mirror descent step to incorporate an optimism vector $o_{t}=o_{t}(r_{1},r_{2},\dots,r_{t-1})$ or $\mu_{t}\cdot\mathbf{1}$, a specific optimism obtained by multiplying a scalar with an all-ones vector, and obtains $p_{t}$, the probability distribution over $N$ bidding prices. This optimism vector can be considered as the learner’s prediction of $r_{t}$. Of course, the closer $o_{t}$ is to $r_{t}$, the smaller static regret rate the learner can achieve. Then the learner chooses $b_{t}$ by sampling from $p_{t}$ and receives $r_{t}$. For the second online mirror descent step, the learner incorporates the actual reward $r_{t}$ to update the knowledge about the environment, possibly with a second-order correction $a_{t}$.
For the case where $L_{T}=o(T)$, we can use Option I of Algorithm 1, which is reminiscent of the earliest instantiation of OMD (Chiang et al. 2012), where $o_{t}=r_{t-1}$. Our choice of optimism follows this idea but we incorporate the information of $v_{t}$ when designing $o_{t}$ to ensure that the private valuation sequence $(v_{t})_{t=1}^{T}$ does not degrade the regret performance. Also, different from the restart scheme in Besbes et al. (2015) using a fixed batch size, we design an adaptive restart procedure to reduce the transition cost from static regret to dynamic regret, and to achieve the minimax-optimal dynamic regret rate.
Input : $\mathcal{P}$ is the convex hull of $\{e_{1},\dots,e_{N}\}$; $\psi(p)$: a convex regularizer defined on the probability simplex.
[…]
for $t\leftarrow 1,\dots,T$ do
   Set
[…]
Choose actions according to $p_{t}$, receive $r_{t,i}$ for any $i\in[N]$ and set
[…]
Update
[…]
end for
1ex
Algorithm 1 Optimistic Mirror Descent
The case of $V_{T}=o(T)$ is more challenging, and we can use Option II of Algorithm 1, which is a variant of OMD by Steinhardt and Liang (2014), Wei and Luo (2018). A key feature of Option II is that the optimism is chosen to be $o_{t}=\mu_{t}\cdot\mathbf{1}$, which might appear rigid at first glance. However, due to this choice and since $p$ lies on the probability simplex, we have $\langle p,\mu_{t}\cdot\mathbf{1}\rangle=\mu_{t}$, which is constant with respect to $p$. Thus, we obtain $p_{t}=p_{t}^{\prime}$ after the first mirror descent step. Therefore, even though $\mu_{t}$ depends on $r_{t}$ $(v_{t}\text{ and }m_{t})$, the variable $p_{t}$ does not depend on $r_{t}$, and we indeed comply with the online learning protocol that requires $p_{t}$ to be chosen before observing $r_{t}$. Our optimism configuration chooses $\mu_{t}=\max\{v_{t}-m_{t},0\}$, which is a novel contribution of our work. The $\mu_{t}$ we choose is not simply targeting the minimization of static regret, but focuses more on achieving a favorable tradeoff between the static regret and the transition cost. This ultimately leads to the minimax-optimal dynamic regret even when using a constant batch size, and we can eliminate the requirement of knowing $V_{T}$ by employing an adaptive batch size.
3.3 Main Results
Our main results are summarized as follows:
Theorem. (informal) For online first-price auctions,
  • consider the set of auction sequences such that $\sum_{t=2}^{T}|m_{t}-m_{t-1}|\leq V_{T}$, then one can apply Algorithm 2 to achieve $\tilde{O}(\sqrt{TV_{T}})$ expected dynamic regret (Theorem 4.6, Section 4.1). Besides, any non-anticipatory policy suffers $\Omega(\sqrt{TV_{T}})$ expected dynamic regret (Theorem 5.2, Section 5.1).
  • consider the set of auction sequences such that $\sum_{t=2}^{T}\mathbbm{1}(m_{t}\neq m_{t-1})\leq L_{T}$, then one can apply Algorithm 3 to achieve $\tilde{O}(L_{T})$ expected dynamic regret (Theorem 4.9, Section 4.2). Besides, any non-anticipatory policy suffers $\Omega(L_{T})$ expected dynamic regret (Theorem 5.4, Section 5.2).
  • consider an auction instance such that $\sum_{t=2}^{T}|m_{t}-m_{t-1}|\leq V_{T}$ and $\sum_{t=2}^{T}\mathbbm{1}(m_{t}\neq m_{t-1})\leq L_{T}$, then Algorithm 4 achieves $\tilde{O}(\sqrt{TV_{T}},L_{T})$ best-of-both-worlds dynamic regret (Theorem 4.10, Section 4.3).
4 Dynamic Regret Upper Bounds
In this section, we explore how to achieve minimax-optimal dynamic regret guarantees under the conditions $V_{T}=o(T)$ or $L_{T}=o(T)$. Our algorithms consist of two main components: a static regret minimizer based on Optimistic Mirror Descent with a carefully chosen optimism vector to handle the one-sided Lipschitzness of the reward function, and a restart scheme with adaptive batch sizes to adapt to the unknown $V_{T}$ without prior knowledge or to achieve an improved dynamic regret rate (in regimes where $L_{T}=o(T)$). Finally, for a specific auction problem instance, it is not a priori clear which regularity metric on the sequence of the opponents’ highest bids leads to a smaller dynamic regret, and we use the meta algorithm by Sani et al. (2014) to establish a best-of-both-worlds dynamic regret guarantee.
4.1 Dynamic Regret Rates under the Temporal Variation Constraint
We first focus on the case where $V_{T}=o(T)$. In Section 4.1.1, we provide a step-by-step illustration of why previous approaches do not work in a straightforward adaptation. In Section 4.1.2, we discuss our minimax-optimal policy, particularly how to design the optimism and how to restart the algorithm with adaptive batch sizes to eliminate the requirement for knowing $V_{T}$ in advance.
4.1.1  Why Existing Works Do Not Directly Apply?
In this section, we discuss several previous approaches that do not work in a straightforward manner for achieving optimal dynamic regret in our setting. These include: (i) the policy proposed by Jadbabaie et al. (2015) for achieving optimal dynamic regret for convex and Lipschitz functions, (ii) restarting the Hedge algorithm with a fixed batch size, and (iii) restarting the policy from Zhang et al. (2022) with a fixed batch size. None of these approaches achieve the optimal dynamic regret rates in our context. Specifically, the approach in Jadbabaie et al. (2015) heavily relies on the Lipschitzness and cannot handle the one-sided Lipschitz reward. The restart Hedge approach fails to adapt to the slowly varying trend of the opponents’ highest bid sequence. While the approach in Zhang et al. (2022) possesses some ability to adapt to the opponents’ highest bid sequence, it lacks sufficient flexibility to optimally balance the static regret and the transition cost, thus leading to a suboptimal dynamic regret rate.
We begin by briefly reviewing the setting of online convex optimization (OCO), as the policy proposed by Jadbabaie et al. (2015) is developed within this framework. OCO models a sequential decision problem as a $T$ round zero-sum game between a learner and an adversary. At round $t$, the learner chooses $x_{t}$ from $\mathcal{X}$, a convex decision set and the adversary reveals $f_{t}$, a convex loss function. An OCO algorithm $\mathcal{A}$ (possibly randomized) maps the historical losses to the current decision: $x_{t}=\mathcal{A}(f_{1},\dots,f_{t-1})\in\mathcal{X}$. The static regret of OCO is defined as
$\mathbb{E}[\text{R}_{T}(\pi)]=\sum_{t=1}^{T}\mathbb{E}[f_{t}(x_{t})]-\min_{x\in\mathcal{X}}\sum_{t=1}^{T}f_{t}(x).$
We refer to $\min_{x\in\mathcal{X}}\sum_{t=1}^{T}f_{t}(x)$ as the static benchmark of OCO. Inspired by non-stationary stochastic optimization problems, Besbes et al. (2015) observe that $\sum_{t=1}^{T}\min_{x_{t}^{*}\in\mathcal{X}}f_{t}(x_{t}^{*})$ (which they term the dynamic benchmark) forms a strictly stronger benchmark. The dynamic regret can be defined as:
$\mathbb{E}[\text{DR}_{T}(\pi)]=\sum_{t=1}^{T}\mathbb{E}[f_{t}(x_{t})]-\sum_{t=1}^{T}\min_{x_{t}^{*}\in\mathcal{X}}f_{t}(x_{t}^{*}).$
[…]
Example 4.1
[…]
We now explore more advanced techniques to address this problem. A key challenge in non-stationary online learning is that the reward sequence may exhibit continuous drifts or abrupt shifts, which diminishes the reliability of older data. Consequently, many existing approaches incorporate mechanisms to “forget” old data, either explicitly or implicitly. In this work, we focus on the restart scheme proposed by Besbes et al. (2015, 2019), partitioning the time horizon $T$ into $n$ batches, denoted by $\mathcal{T}_{j}$, each of length $\Delta_{T,j}$. While Besbes et al. (2015) consider fixed batch lengths, we allow varying lengths for greater flexibility. Adapting Besbes et al. (2015, Proposition 2) to our online first-price auction problem, the dynamic regret can be decomposed as follows:
[…]
We decompose the dynamic regret over the time horizon $T$ into contributions from $n$ batches. The dynamic regret within each batch $\mathcal{T}_{j}$ is further decomposed into the static regret and a transition cost. Specifically, $\mathcal{S}^{\mathcal{A}}(\tilde{\mathcal{F}},\mathcal{T}_{j})$ denotes the static regret of algorithm $\mathcal{A}$ applied to batch $\mathcal{T}_{j}$ against the best fixed policy in a policy class $\tilde{\mathcal{F}}$. The term $\mathcal{C}(\tilde{\mathcal{F}},\mathcal{T}_{j})$ represents the transition cost from static to dynamic regret for batch $\mathcal{T}_{j}$ and policy set $\tilde{\mathcal{F}}$.
To demonstrate the application of the decomposition in Equation (6) for achieving sublinear dynamic regret, we consider using the restart scheme with the Hedge algorithm as $\mathcal{A}$. We partition the time horizon into batches of equal length $\Delta_{T}$, with the possible exception of the last batch. Let $\tilde{\mathcal{F}}$ be the set of constant policies, i.e., $\tilde{\mathcal{F}}\coloneqq\{f(v;\tau)=\tau|\tau\in[0,1]\}$. Then, the following proposition holds:
Proposition 4.2
Assume $V_{T}=o(T)$ and is known, $V_{T}^{v}\coloneqq\sum_{t=2}^{T}|v_{t}-v_{t-1}|=o(T)$, then restarting the Hedge policy every $\Delta_{T}$ rounds, where $\Delta_{T}=O\left(\left(\frac{T}{V_{T}+V_{T}^{v}}\right)^{\frac{2}{3}}\right)$ achieves $\tilde{O}\left(T^{\frac{2}{3}}(V_{T}+V_{T}^{v})^{\frac{1}{3}}\right)$ dynamic regret.
Finally, we examine the results of Zhang et al. (2022), which study online first-price auctions where a hint $h_{t}$ is provided before deciding the bid $b_{t}$. The hint satisfies $\mathbb{E}\left[|h_{t}-m_{t}|^{q}\right]\leq\sigma_{t}^{q}$ for any $t\in[T]$. The single hint setting is considered in their work, where they assume an upper bound on $\sum_{t=1}^{T}\sigma_{t}$ is available. Our problem can be viewed as a special case of the single hint setting by choosing $h_{t}=m_{t-1}$, $q=1$, and $V_{T}=\sum_{t=1}^{T}\sigma_{t}$. Then, Zhang et al. (2022, Theorem 2) demonstrate that, when $v_{t}\equiv 1$, there exists an algorithm that guarantees the following static regret bound:
$\mathbb{E}[\text{R}_{T}(\pi)]=\max_{f\in\mathcal{F}_{\text{Lip}}}\sum_{t=1}^{T}r(f(v_{t});v_{t},m_{t})-\sum_{t=1}^{T}\mathbb{E}\left[r(b_{t};v_{t},m_{t})\right]=\tilde{O}\left(T^{\frac{1}{4}}V_{T}^{\frac{1}{4}}\right),$ (7)
where $\mathcal{F}_{\text{Lip}}$ is the set of $1$-Lipschitz policies $f:[0,1]\to[0,1]$.
By combining the restart scheme with the algorithm in Zhang et al. (2022, Theorem 2), we obtain the following result.
Proposition 4.3
Assume $v_{t}\equiv 1$ holds for $t\in[T]$ and $V_{T}=o(T)$ is known, then the learner can restart the algorithm in Zhang et al. (2022, Theorem 2) every $\Delta_{T}$ rounds, where $\Delta_{T}=O\left(\left(\frac{T}{V_{T}}\right)^{\frac{1}{2}}\right)$ to achieve $\tilde{O}\left(\sqrt{TV_{T}}\right)$ dynamic regret.
Proposition 4.2 establishes an $\tilde{O}(T^{2/3}(V_{T}+V_{T}^{v})^{1/3})$ upper bound on the dynamic regret. However, this bound is suboptimal compared to the $\Omega(\sqrt{TV_{T}})$ lower bound presented in Theorem 5.2. While Proposition 4.3 achieves the optimal rate, it relies on the restrictive assumption that $v_{t}\equiv 1$ for $t\in[T]$. Although Zhang et al. (2022) consider varying $v_{t}$ as well, they employ the ChEW policy (Han et al. 2020), which is an inefficient variant of the SEW policy (also from Han et al. (2020)), to achieve $\tilde{O}(\sqrt{T})$ static regret. Directly combining this rate with the restart scheme and following the proof of Proposition 4.3 results in a dynamic regret of $\tilde{O}\left(T^{\frac{2}{3}}V_{T}^{\frac{1}{3}}\right)$, which is still suboptimal.
Consequently, achieving the optimal dynamic regret rate for varying $v_{t}$ using existing approaches remains an open problem. We will explore alternative methodologies to address this.
4.1.2  Minimax-Optimal Policy and Parameter-free Scheme.
In this section, we investigate how to achieve the minimax-optimal dynamic regret upper bound. Our main approach is to design a suitable restart scheme that employs the framework of Optimistic Mirror Descent (Algorithm 1, Option II) as the static regret minimizer. The key technical contribution is to provide a novel optimism configuration $\mu_{t}=\max\{v_{t}-m_{t},0\}$, which yields a favorable balance between the static regret and the transition cost, thus leading to the optimal dynamic regret rate. When $V_{T}$ is known, a constant batch size suffices to achieve the optimal dynamic regret. When $V_{T}$ is unknown, we can employ an adaptive batch size (Algorithm 2) to achieve the optimal dynamic regret as well. We provide insights and details about the optimal policy below.
Recall that the proof in Proposition 4.2 is based on the following argument:
$\begin{split}\mathbb{E}[\text{DR}_{T}(\pi)]&=\sum_{j=1}^{n}\mathcal{S}^{\mathcal{A}}(\tilde{\mathcal{F}},\mathcal{T}_{j})+\sum_{j=1}^{n}\mathcal{C}(\tilde{\mathcal{F}},\mathcal{T}_{j})\\ &\leq\left\lceil\frac{T}{\Delta_{T}}\right\rceil\cdot\tilde{O}\left(\sqrt{\Delta_{T}}\right)+\Delta_{T}(V_{T}+V_{T}^{v})\\ &=\tilde{O}\left(\frac{T}{\sqrt{\Delta_{T}}}+\Delta_{T}(V_{T}+V_{T}^{v})\right)=\tilde{O}\left(T^{\frac{2}{3}}(V_{T}+V_{T}^{v})^{\frac{1}{3}}\right)\end{split}$ (8)
with optimal tuning of the batch size $\Delta_{T}$. The $O(\sqrt{\Delta_{T}})$ static regret achieved with this tuning, while minimax-optimal for each batch $j\in[n]$, is not tight when batch $j$’s temporal variation, $V_{T,j}=\sum_{t\in\mathcal{T}_{j}}|m_{t}-m_{t-1}|$, is significantly smaller than $\Delta_{T}$. For instance, if the values $m_{t}$ are constant within batch $\mathcal{T}_{j}$, we expect $O(1)$ static regret rather than the minimax-optimal $O(\sqrt{\Delta_{T}})$. This observation leads us to investigate the existence of online learning policies with static regret bounds that scale with the temporal variation of the opponents’ highest bid sequence $(m_{t})_{t=1}^{T}$. In the machine learning theory community, this question aligns with the concept of adaptive online learning (Cesa-Bianchi et al. 2007, Rakhlin and Sridharan 2013, Wei and Luo 2018), which focuses on achieving static regret guarantees that scale with the “complexity” of the input data.
Inspired by this observation and the $\Omega(\sqrt{TV_{T}})$ lower bound that we will establish in Section 5.1, we conjecture that an improved dynamic regret bound can be achieved by considering:
$\begin{split}\mathbb{E}[\text{DR}_{T}(\pi)]&=\sum_{j=1}^{n}\mathcal{S}^{\mathcal{A}^{\prime}}(\mathcal{F}^{\prime},\mathcal{T}_{j})+\sum_{j=1}^{n}\mathcal{C}(\mathcal{F}^{\prime},\mathcal{T}_{j})\\ &\stackrel{{\scriptstyle?}}{{\leq}}\sum_{j=1}^{\lceil{T/\Delta_{T}}\rceil}\tilde{O}\left(\Delta_{T}V_{T,j}+1\right)+\Delta_{T}V_{T}\\ &=\tilde{O}\left(\Delta_{T}V_{T}+\frac{T}{\Delta_{T}}\right)+\Delta_{T}V_{T}=\tilde{O}\left(\Delta_{T}V_{T}+\frac{T}{\Delta_{T}}\right)=\tilde{O}\left(\sqrt{TV_{T}}\right),\end{split}$ (9)
[…]
While it may initially seem surprising that such an adaptive policy could improve dynamic regret, given that $\Delta_{T}V_{T,j}$ can exceed $\sqrt{\Delta_{T}}$ for some $j\in[n]$, the adaptive nature of $\mathcal{A}^{\prime}$ and the fact that $\sum_{j=1}^{n}V_{T,j}\leq V_{T}$ allow for a more favorable balance between the overall static regret $\sum_{j=1}^{n}\mathcal{S}^{\mathcal{A}^{\prime}}(\mathcal{F}^{\prime},\mathcal{T}_{j})$ and the overall transition cost $\sum_{j=1}^{n}\mathcal{C}(\mathcal{F}^{\prime},\mathcal{T}_{j})$. This permits a more aggressive choice of $\Delta_{T}$, leading to an improved dynamic regret rate. We refer to this idea as “adaptive balancing,” as it leverages adaptive online learning algorithms to balance the scales of the static regret and the transition cost.
To achieve an $\tilde{O}(\Delta_{T}V_{T,j}+1)$ static regret bound, we require an algorithm satisfying two conditions: (i) its regret should scale with the temporal variation of the sequence $(m_{t})_{t\in\mathcal{T}_{j}}$, and (ii) it should be customizable to facilitate adaptive balancing. The OMD framework (Chiang et al. 2012, Rakhlin and Sridharan 2013) fulfills both requirements. In particular, we focus on Option II of Algorithm 1, which implies a static regret bound of the form $O\left(\sqrt{\sum_{t=1}^{T}(r_{t,i^{*}}-\mu_{t})^{2}\ln N}\right)$ (Steinhardt and Liang 2014, Wei and Luo 2018), where $i^{*}$ is the index of the optimal expert in hindsight. As mentioned in Section 3.2, the optimism vector $\mu_{t}\cdot\mathbf{1}$ plays an important role in balancing the static regret and the transition cost. We choose
$\mu_{t}=\max\{v_{t}-m_{t},0\},$ (10)
which is a novel contribution of this work. We provide intuition about how we derive this optimism below. Notably, this choice of $\mu_{t}$ coincides with $r(b_{t}^{*};v_{t},m_{t})$ in Equation (1), which helps to relate the static regret and the transition cost—a point that will be more transparent in the proof of Theorem 4.6.
As discussed in Section 3.2, when the optimism vector is a constant times an all-ones vector, such as $\mu_{t}\cdot\mathbf{1}$, $\mu_{t}$ can depend on $r_{t}$, the reward at round $t$. Since the reward $r_{t}$ depends on both the private valuation $v_{t}$ and the opponents’ highest bid $m_{t}$, it is natural to parametrize $\mu_{t}$ as a function of these two variables. We assume $\mu_{t}=\mu(v_{t},m_{t})$, which turns out to make our theory work after some calculations. Next, we discuss how to determine the optimism $\mu_{t}=\mu(v_{t},m_{t})$.
When we restrict our focus to $\mathcal{T}_{j}$, the $j$-th batch, we need an algorithm with regret upper bounded by $\tilde{O}(1+\Delta_{T}V_{T,j})$, as illustrated in Equation (9). We analyze the problem instance in Example 4.4, which contains several parameters like $v,m$ and $\tilde{m}$. By examining different regimes of these parameters, we find that choosing $\mu_{t}$ as in Equation (10) is indeed reasonable. This optimism can be combined with the fact that $i^{*}$ is the optimal expert to show the desired adaptive static regret bound. While here we gain insights using special examples, later we find this optimism indeed works in general. Therefore, we can achieve $\tilde{O}(1+\Delta_{T}V_{T,j})$ static regret by combining Option II of Algorithm 1 with Equation (10).
Example 4.4
Consider the following first-price auction instance on batch $\mathcal{T}_{j}$, $v_{t}\equiv v$ for $t\in\mathcal{T}_{j}$ and
$(m_{t})_{t\in\mathcal{T}_{j}}=(\underbrace{m,m,\dots,m}_{T_{1}\text{ copies }},\underbrace{\tilde{m},\tilde{m},\dots,\tilde{m}}_{T_{2}\text{ copies}}),$
where $T_{1}+T_{2}=\Delta_{T}$, the batch size.
However, computing $p_{t+1}^{\prime}$ in Algorithm 1 with Option II requires solving a convex optimization problem, which can be computationally expensive. Therefore, we employ the Prod forecaster (Cesa-Bianchi et al. 2007), which offers the same $O(\sqrt{\sum_{t=1}^{T}(r_{t,i^{*}}-\mu_{t})^{2}\ln N})$ regret guarantee with more efficient updates:
$p_{1}=\left(\frac{1}{N},\dots,\frac{1}{N}\right),\qquad p_{t+1,i}=\frac{(1+\eta(r_{t,i}-\mu_{t}))p_{t,i}}{\sum_{j=1}^{N}(1+\eta(r_{t,j}-\mu_{t}))p_{t,j}},$ (11)
Furthermore, the dynamic regret bound in Proposition 4.2 has an undesirable dependence on $V_{T}^{v}\coloneqq\sum_{t=2}^{T}|v_{t}-v_{t-1}|$. We aim to eliminate this dependence, which arises from the one-sided Lipschitz property of the reward function:
Lemma 4.5
(Han et al. 2020) For any $v,m\in[0,1]$, $b\leq\min\{v,b^{\prime}\}$,
$r(b;v,m)-r(b^{\prime};v,m)\leq b^{\prime}-b.$
Lemma 4.5 implies that the one-sided Lipschitzness of the reward function relies on the condition $b\leq v$, meaning that the set of constant policies does not satisfy this property. Notably, Han et al. (2020) encountered a similar difficulty, where they aimed to compete with the best fixed policy within the set of $1$-Lipschitz policies $\mathcal{F}_{\text{Lip}}$. However, they found that restricting the policy set to $\mathcal{F}_{0}\coloneqq\{f\mid f\in\mathcal{F}_{\text{Lip}},f(v)\leq v\}$ does not compromise the reward and resolves the problem. Inspired by this, we define $\mathcal{F}\coloneqq\{f(v;\tau)\mid\tau\in[0,1]\}$, where $f(v;\tau)\coloneqq\min\{v,\tau\}$. $\mathcal{F}$ can be viewed as a modified version of $\mathcal{N}\coloneqq\{\tau\mid\tau\in[0,1]\}$, the set of constant policies, with the additional constraint $f(v;\tau)\leq v$. We further define $\mathcal{F}_{\epsilon}\coloneqq\{f(v;\tau)\mid\tau\in\{0,\epsilon,2\epsilon,\dots,\epsilon\lfloor 1/\epsilon\rfloor\}\}$, which is a discretized version of $\mathcal{F}$ with precision $\epsilon$. Using this setup, we can effectively eliminate the dependence on $V_{T}^{v}$ through a careful application of the one-sided Lipschitzness property given in Lemma 4.5.
With all the necessary tools in place, we now illustrate how to leverage the concept of “adaptive balancing” to achieve an improved dynamic regret rate. Assuming $V_{T}$ is known, it is sufficient to restart the Prod forecaster every $\Delta_{T}$ rounds, where $\Delta_{T}=O\left(\sqrt{\frac{T}{V_{T}}}\right)$, to achieve a dynamic regret of $\tilde{O}\left(\sqrt{TV_{T}}\right)$. However, in practice, $V_{T}$ is typically unknown. To address this, we use an adaptive restart condition, as demonstrated in Algorithm 2, to resolve the issue while still achieving the minimax-optimal rate. Theorem 4.6 establishes the minimax-optimal dynamic regret guarantee under the condition $V_{T}=o(T)$.
Input : Time horizon $T$
[…]
while $t\leq T$ do
   Observe the ad impression at $t$ and generate the value $v_{t}$;
   Create $\mathcal{T}_{j}$;
 ++j++j;
 $p_{t}\leftarrow\left(\frac{1}{N},\dots,\frac{1}{N}\right)$ where $N\leftarrow\frac{1}{\epsilon}$;
 while $\Delta_{T,j}<\sqrt{\frac{T}{\sum_{i=1}^{j}V_{T,i}+c}}$ do
      Choose $b_{t}\leftarrow\min\{v_{t},i\epsilon\}$ with probability $p_{t,i}$;
      Submit $b_{t}$ and receive $m_{t}$;
      Update $\Delta_{T,j}$ and $V_{T,j}$;
    // $\Delta_{T,j}$: length of $\mathcal{T}_{j}$, $V_{T,j}$: temporal variation of $(m_{t})_{t\in\mathcal{T}_{j}}$
    $\mu_{t}\leftarrow\max\{v_{t}-m_{t},0\}$;
    $p_{t+1,i}\leftarrow\frac{(1+\eta(r_{t,i}-\mu_{t}))p_{t,i}}{\sum_{j=1}^{N}(1+\eta(r_{t,j}-\mu_{t}))p_{t,j}}$;
      ++t;
      Observe the ad impression at $t$ and generate the value $v_{t}$;
   end while
end while
Algorithm 2 The Adaptive Restart Prod Policy (AR-Prod)
Theorem 4.6
Assume $V_{T}=o(T)$. When $V_{T}$ is known, we can restart the Prod forecaster (Equation (11)) with $\mu_{t}=\max\{v_{t}-m_{t},0\}$ using a constant batch size $\Delta_{T}=O\left(\sqrt{\frac{T}{V_{T}}}\right)$ to achieve the $\tilde{O}\left(\sqrt{TV_{T}}\right)$ dynamic regret. When $V_{T}$ is unknown, Algorithm 2 restarts the Prod policy adaptively and achieves
$\sup_{(v_{t},m_{t})_{t=1}^{T}\in\mathcal{V}}\mathbb{E}\left[\textnormal{DR}_{T}(\pi)\right]=\tilde{O}\left(\max\left\{\sqrt{TV_{T}},1\right\}\right).$
Proof 4.7
Proof Sketch. We begin by considering the case where $V_{T}$ is known. The proof follows the approach suggested in Equation (9). We define $\mathcal{F}\coloneqq\{f(v;\tau)\mid\tau\in[0,1]\}$ and $\mathcal{F}_{\epsilon}\coloneqq\left\{f(v;\tau)\mid\tau=k\epsilon,k\in\left[\left\lfloor\frac{1}{\epsilon}\right\rfloor\right]\right\}$ as the set of policies and its discretization, respectively, where $f(v;\tau)\coloneqq\min\{v,\tau\}$. We first consider the case where $V_{T}$ is known. We divide the time horizon $T$ into batches $\mathcal{T}_{1},\mathcal{T}_{2},\dots,\mathcal{T}_{n}$ of equal length (possibly except $\mathcal{T}_{n}$) and consider the dynamic regret
$\begin{split}\mathbb{E}[\text{DR}_{T}(\pi)]&=\sum_{j=1}^{n}\mathcal{S}^{\mathcal{A}}(\mathcal{F},\mathcal{T}_{j})+\sum_{j=1}^{n}\mathcal{C}(\mathcal{F},\mathcal{T}_{j})\\ &\leq\sum_{j=1}^{\lceil{T/\Delta_{T}}\rceil}\tilde{O}\left(\Delta_{T}V_{T,j}+1\right)+\sum_{j=1}^{\lceil{T/\Delta_{T}}\rceil}\Delta_{T}V_{T,j}\\ &=\tilde{O}\left(\Delta_{T}V_{T}+\frac{T}{\Delta_{T}}\right)+\Delta_{T}V_{T}=\tilde{O}\left(\Delta_{T}V_{T}+\frac{T}{\Delta_{T}}\right)=\tilde{O}\left(\sqrt{TV_{T}}\right),\end{split}$ (12)
where $\mathcal{A}$ is the Prod forecaster illustrated in Equation (11). The inequality is shown by the following idea: by choosing the translation term in the Prod forecaster as $\mu_{t}\coloneqq\max\{v_{t}-m_{t},0\}$, we can show
$S^{\mathcal{A}}(\mathcal{F},\mathcal{T}_{j})=O(\mathcal{C}(\mathcal{F},\mathcal{T}_{j}))+\tilde{O}(1)$ (13)
holds for any batch $j$. It then suffices to show $\mathcal{C}(\mathcal{F},\mathcal{T}_{j})\leq\Delta_{T}V_{T,j}$ to establish that the inequality in Equation (13) holds, which is possible since our expert set $\mathcal{F}$ facilitates the application of Lemma 4.5.
Now suppose $V_{T}$ is unknown, then we use the adaptive restart routine in Algorithm 2. Following the argument in Equation (12), we can establish
$\mathbb{E}[\text{DR}_{T}(\pi)]=\tilde{O}\left(n+\sum_{j=1}^{n}\Delta_{T,j}V_{T,j}\right),$ (14)
where $n$ denotes the number of batches, $\Delta_{T,j}$ represents the length of batch $j$, and $V_{T,j}$ denotes the temporal variation of $m_{t}$ within batch $j$. While these quantities ($n$, $\Delta_{T,j}$, and $V_{T,j}$) are a priori unknown, leveraging the restart condition in conjunction with the self-confident tuning technique (cf. Auer et al. (2002)) allows us to effectively bound them. Specifically, these techniques yield $n=\tilde{O}(\sqrt{TV_{T}})$ and $\sum_{j=1}^{n}\Delta_{T,j}V_{T,j}=\tilde{O}(\sqrt{TV_{T}})$, where $T$ is the total time horizon and $V_{T}$ denotes the total temporal variation across all batches. Consequently, substituting these bounds into Equation (14) yields the desired $\tilde{O}(\sqrt{TV_{T}})$ bound.
Remark 4.8
Previous proofs for learning in non-stationary environments (Besbes et al. 2015, 2019, Cheung et al. 2022, 2023) typically decompose the dynamic regret into the sum of static regret and transition cost, and then bound these terms individually. While this approach could also be applied to our problem, in the proof of Theorem 4.6, we instead establish a direct relationship between the static regret and the transition cost (Equation (13)). This alternative approach results in a more transparent proof and may be of independent interest.
4.2 Dynamic Regret Rates under the Switching Number Constraint
We now consider the case where the number of switches in the opponents’ highest bid sequence, $L_{T}=\sum_{t=2}^{T}\mathbbm{1}(m_{t}\neq m_{t-1})$, is $o(T)$. Our approach combines the Optimistic Mirror Descent (OMD) framework (Algorithm 1, Option I) with an adaptive restart scheme: OMD with a suitable optimism vector $o_{t}$ is run within each batch, and a new batch is started whenever a change in $m_{t}$ is detected (i.e., $m_{t}\neq m_{t-1}$).
Since $m_{t}$ is observed exactly, each batch contains at most one switch ($m_{t}\neq m_{t-1}$). Due to the configured optimism, the static regret for each batch corresponds to the number of switches. Given this single-switch property, we can show the transition cost from static regret to dynamic regret is $\tilde{O}(1)$. Combining both parts, and summing over all $L_{T}$ batches, the total dynamic regret is $\tilde{O}(L_{T})$. We use the negative entropy regularizer in OMD, which allows for efficient closed-form updates as shown in Algorithm 3. Theorem 4.9 formalizes this result, establishing a dynamic regret upper bound.
Input : $\mathcal{P}$ is the convex hull of $\{e_{1},\dots,e_{N}\}$; $\psi(p)\leftarrow\frac{1}{\eta}\sum_{i=1}^{N}p_{i}\ln p_{i}$
$j\leftarrow 1$$t\leftarrow 1$;
1exwhile $t\leq T$ do
   Create $\mathcal{T}_{j}$;
   ++ $j$;
   Update
$p_{t,i}\propto\exp\left(\eta\left(\sum_{s=1}^{t-1}r_{s,i}+o_{t,i}\right)\right)$
where $o_{t,i}\coloneqq r(f(v_{t};i\epsilon);v_{t},m_{t-1})$;
   Submit bids according to $p_{t}$, and receive $m_{t}$;
 while $t$ is the first round in $\mathcal{T}_{j}$ or $m_{t}=m_{t-1}$ do
    ++t++t;
      Update
$p_{t,i}\propto\exp\left(\eta\left(\sum_{s=1}^{t-1}r_{s,i}+o_{t,i}\right)\right)$
where $o_{t,i}\coloneqq r(f(v_{t};i\epsilon);v_{t},m_{t-1})$;
   end while
end while
1ex
Algorithm 3 Adaptive Restart Optimistic Mirror Descent (AR-OMD)
Theorem 4.9
Assume $L_{T}=o(T)$ and is unknown, then Algorithm 3 achieves
$\sup_{(v_{t},m_{t})_{t=1}^{T}\in\mathcal{L}}\mathbb{E}\left[\textnormal{DR}_{T}(\pi)\right]=\tilde{O}\left(L_{T}\right).$
4.3 Best-of-Both-Worlds Dynamic Regret
In Sections 4.1 and 4.2, we have established the $\tilde{O}(\sqrt{V_{T}T})$ and $\tilde{O}(L_{T})$ dynamic regret rates for slowly varying and abruptly changing bidding environments, respectively. But in reality, it is hard for a learner to know a priori which non-stationary measure is suitable, thus it is desirable to automatically achieve the better of the two guarantees whenever one outperforms the other. This problem is termed as the best-of-both-worlds bound in the online learning literature. An important technique for establishing the best-of-both-worlds bound is to run a few base algorithms in parallel, and use a meta algorithm to aggregate the output of base algorithms to ensure the resulting long-term performance is always as good as the best base algorithm. In this part, we establish the best-of-both-worlds bound based on the meta algorithm by Sani et al. (2014). The resulting algorithm and theoretical guarantee are presented as Algorithm 4 and Theorem 4.10, respectively, and the proof of Theorem 4.10 is deferred to Appendix 4.3.
Theorem 4.10
Assume $V_{T}=\Omega(\ln T)$, then Algorithm 4 achieves $\tilde{O}(\min\{\sqrt{TV_{T}},L_{T}\})$ best-of-both-worlds dynamic regret guarantee for online first-price auctions.
Input : Let $\mathcal{A}$ and $\mathcal{B}$ be Algorithms 2 and 3, respectively; total number of rounds $T$, learning rate $\eta=\frac{1}{2}\cdot\sqrt{\frac{\ln T}{T}}$, initial weights $w_{1}^{\mathcal{A}}=\eta,w_{1}^{\mathcal{B}}=1-\eta$.
$j\leftarrow 1$$t\leftarrow 1$;
1exfor $t\leftarrow 1$ to $T$ do
 $p_{t}=\frac{w_{t}^{\mathcal{A}}}{w_{t}^{\mathcal{A}}+w_{t}^{\mathcal{B}}}$;
   Observe bids $b_{t}^{\mathcal{A}}$ and $b_{t}^{\mathcal{B}}$ produced by $\mathcal{A}$ and $\mathcal{B}$;
   Bid
$b_{t}=\begin{cases}b_{t}^{\mathcal{A}},&\text{with probability }p_{t},\\ b_{t}^{\mathcal{B}},&\text{otherwise},\end{cases}$
Observe $m_{t}$ and get reward $r(b_{t};v_{t},m_{t})$;
   Send $m_{t}$ to $\mathcal{A}$ and $\mathcal{B}$;
   Let $\delta_{t}=r_{t}(b_{t}^{\mathcal{A}};v_{t},m_{t})-r_{t}(b_{t}^{\mathcal{B}};v_{t},m_{t})$;
   Set $w_{t+1}^{\mathcal{A}}=w_{t}^{\mathcal{A}}(1+\eta\delta_{t})$;
end for
1ex
Algorithm 4 Non-stationary First-price Auction with Best-of-Both-Worlds Guarantee.
Remark 4.11
If we want to obtain a best-of-both-worlds static regret bound for two algorithms with different adaptive static regret guarantees, the meta algorithm by Sani et al. (2014) might not be applicable since the overhead can be as large as $O(\sqrt{T\ln T})$ for one adaptive regret guarantee, while it is $O(1)$ for the other adaptive regret guarantee. The $O(\sqrt{T\ln T})$ overhead can destroy an adaptive regret guarantee. Fortunately, the dynamic regret guarantee in Theorem 4.6 is $\tilde{O}(\sqrt{TV_{T}})$, which can easily absorb the $O(\sqrt{T\ln T})$ overhead as long as $V_{T}=\Omega(\ln T)$.
5 Dynamic Regret Lower Bounds
In this section, we demonstrate how to establish minimax lower bounds for the class of auction instances where the opponents’ highest bid sequence is constrained by either $V_{T}$ or $L_{T}$. Our main effort focuses on the case of $V_{T}$. Following Besbes et al. (2015), we term $V_{T}$ the variation budget. For the corresponding lower bound construction, we partition the time horizon $T$ into batches of equal size $H$, and we allocate a small and fixed amount of variation budget $\frac{1}{H}$ to each batch to create a jump at locations drawn from the uniform distribution. Within each batch, due to the one-sided Lipschitzness of the reward function, the learner faces a dilemma: providing a small bid incurs 0 dynamic regret before the jump occurs, but will incur $\Omega(1)$ dynamic regret at the jump point. Bidding a higher price avoids the $\Omega(1)$ dynamic regret at the jump point, but incurs $\frac{1}{H}$ dynamic regret for each round until the jump point. Formally, we show that any non-anticipatory policy incurs $\Omega(1)$ dynamic regret within each batch based on dynamic programming in Lemma 5.1. Since there are $\Theta(T/H)$ batches, choosing $H=\Theta(\sqrt{T/V_{T}})$ satisfies the variation budget constraint and also implies the $\Omega(\sqrt{TV_{T}})$ dynamic regret lower bound. For the case of $L_{T}$, we achieve the $\Omega(L_{T})$ lower bound by reducing to the case of $V_{T}$, which is possible because each batch considered in Lemma 5.1 contains only one random jump.
5.1 Minimax Lower Bound under the Temporal Variation Constraint.
In this section, we establish an $\Omega(\sqrt{TV_{T}})$ lower bound for online first-price auctions. We begin by outlining the technical challenges. The adversary’s objective is to optimally allocate the variation budget across the entire time horizon. Existing lower bounds for learning in non-stationary environments (Besbes et al. 2015, 2019, Cheung et al. 2022) typically rely on the presence of noisy feedback. This noise allows the construction of two reward functions and the partitioning of the time horizon into batches of size $\Delta_{T}$. Within each batch, a reward function is selected uniformly at random and applied consistently. The noisy feedback ensures that, within each batch, the learner perceives i.i.d. rewards. Consequently, information-theoretic tools can be employed to lower bound the probability of identifying the true underlying reward function. However, in our setting, the learner observes $m_{t}$ directly, without noise. This absence of noise necessitates the development of alternative approaches.
In this work, for the lower bound construction, we design problem instances that satisfy the variation budget constraint, where the dynamic regret of any non-anticipatory policy can be computed using dynamic programming. Similar ideas have been used to derive minimax lower bounds on the static regret for learning with a small number of experts (Cover 1966, Gravin et al. 2016, Harvey et al. 2023). Specifically, our approach constructs an opponents’ highest bid sequence of length $H$ with temporal variation bounded by $1/H$, showing that any admissible policy incurs $\Omega(1)$ dynamic regret on this sequence. By concatenating $\Theta(T/H)$ such sequences with $H=\Theta\left(\sqrt{T/V_{T}}\right)$, we construct a total sequence with temporal variation bounded by $V_{T}$. The total dynamic regret is then lower bounded by the number of sequences multiplied by $\Omega(1)$, yielding $\Omega(T/H)=\Omega(\sqrt{TV_{T}})$, as desired.
Lemma 5.1 provides the construction of a single sequence and establishes the $\Omega(1)$ lower bound on its dynamic regret.
Lemma 5.1
Let $H\geq 2$ be an integer, and consider a $H$-round online first-price auction game, assume $v_{t}\equiv 1$, and
$m_{t}=\begin{cases}0,&t<\tau,\\ \delta,&\tau\leq t\leq H,\end{cases}$
where $\tau$ is uniformly drawn from $\{1,2,\dots,H\}$. Then any non-anticipatory policy suffers at least $\frac{1}{2}-\frac{1}{2H}$ dynamic regret when $\delta=\frac{1}{H}$.
Lemma 5.1 shows that within a batch of length $H$, a variation budget of $1/H$ can induce $\Omega(1)$ dynamic regret for any admissible algorithm. With a total variation budget of $V_{T}$, we can construct $\Theta(HV_{T})$ such batches. To achieve the desired lower bound, we set $H=\sqrt{T/V_{T}}$. However, directly concatenating $T/H$ batches as described in Lemma 5.1 would result in $m_{t}$ reaching $\frac{1}{H}\cdot\frac{T}{H}=V_{T}$ at later stages, potentially violating the assumption that $m_{t}\in[0,1]$.
To address this issue, we employ an alternating batch construction. We divide the time horizon into batches of length $H$ and indexed by $j$. For odd $j$, we use the batches constructed in Lemma 5.1. For even $j$, we use batches defined as follows:
$m_{t}=\begin{cases}\delta,&t<\tau,\\ 0,&\tau\leq t\leq H,\end{cases}$
where $\tau$ is drawn uniformly from $1,2,\dots,H$} and $\delta=\frac{1}{H}$. This alternating construction ensures that $m_{t}\in[0,1]$, as depicted in Figure 1. By alternating between these types of batches, we can fully utilize the variation budget while respecting the constraint that $m_{t}\in[0,1]$.
Figure 1: An illustration of the construction for the lower bound. In odd-numbered batches, $m_{t}$ jumps from 0 to $\frac{1}{H}$ at random locations, while in even-numbered batches, $m_{t}$ jumps from $\frac{1}{H}$ back to 0. The parameter $H$ is carefully chosen to satisfy the variation budget constraint and ensure that we obtain the desired lower bound.
We stitch batches constructed in Lemma 5.1, and establish the minimax-optimal lower bound as follows:
Theorem 5.2
In online first-price auctions, for any $V_{T}\in\left[\frac{36}{T},\frac{T}{4}\right]$, there exists $(v_{t},m_{t})_{t=1}^{T}$ such that $\sum_{t=2}^{T}|m_{t}-m_{t-1}|\leq V_{T}$ and the expected dynamic regret of any admissible policy satisfies:
$\inf_{\pi\in\Pi}\sup_{(v_{t},m_{t})_{t=1}^{T}\in\mathcal{V}}\mathbb{E}\left[\textnormal{DR}_{T}(\pi)\right]\geq\frac{\sqrt{TV_{T}}}{16},$
where $\Pi$ is the set of admissible policies.
Remark 5.3
Due to the construction of the lower bound, we can explicitly inform the learner about the creation of the opponents’ highest bid batches, the variation budget allocated to each batch, and the total number of batches. The lower bound remains valid under this setting. This implies that our lower bound holds even when the learner is aware of $V_{T}$, whereas our upper bound does not require prior knowledge of $V_{T}$.
5.2 Minimax Lower Bound under the Discrete Switching Constraint.
We also establish a corresponding minimax lower bound for the case of $L_{T}=o(T)$ by reducing it to the proof of Theorem 5.2.
Theorem 5.4
In online first-price auctions, for any $L_{T}\in[T]$ and $L_{T}\leq\frac{T}{3}$, there exists $(v_{t},m_{t})_{t=1}^{T}$ such that $\sum_{t=2}^{T}\mathbbm{1}(m_{t}\neq m_{t-1})\leq L_{T}$ and the expected dynamic regret of any admissible policy satisfies:
$\inf_{\pi\in\Pi}\sup_{(v_{t},m_{t})_{t=1}^{T}\in\mathcal{L}}\mathbb{E}\left[\textnormal{DR}_{T}(\pi)\right]\geq\frac{L_{T}}{8}.$
It is insightful to compare Theorems 5.2 and 5.4 with Proposition 3.6. Proposition 3.6 essentially establishes an $\Omega(V_{T})$ lower bound for $V_{T}\in\left[0,\frac{T}{2}\right]$ and an $\Omega(L_{T})$ lower bound when $L_{T}=\Theta(T)$. In contrast, the lower bounds in Theorems 5.2 and 5.4 are significantly stronger.
6 Numerical Experiments
In this section, we conduct numerical experiments to evaluate the performance of our proposed algorithms and compare them with baseline methods. Our experiments consist of two main parts: in the first part, we generate the sequence of opponents’ highest bids based on four slowly varying patterns as considered in Besbes et al. (2015, 2019), Cheung et al. (2022), and then confirm our theoretical findings by evaluating the slopes of the log-log plots of dynamic regret with respect to different time horizons. In the second part, we run our proposed algorithms as well as two baseline methods to bid in a multi-agent bidding environment, where each opponent applies the budget-pacing policy from Gaitonde et al. (2022). We find that our algorithms outperform the baselines, especially in regimes where opponents have limited budgets, even when the sequence of opponents’ values varies rapidly.
6.1 Dynamic Regret Growth with Varying Time Horizons
We consider four different patterns for the opponents’ highest bid sequence $(m_{t})_{t=1}^{T}$ and then evaluate the slopes of the log-log plots of dynamic regret with respect to different time horizons to confirm the rates predicted by our theoretical findings. The first three patterns are constructed using the following building blocks from Besbes et al. (2015):
$m_{t}^{\text{constant}}=\begin{cases}0,&t\leq\tau,\\ 1,&t>\tau\end{cases}\quad m_{t}^{\text{exponential}}=\begin{cases}0,&t\leq\tau,\\ 1-e^{-10(t-\tau)/T},&t>\tau\end{cases}\quad m_{t}^{\text{linear}}=\begin{cases}0,&t\leq\tau,\\ \frac{t-\tau}{T-\tau},&t>\tau\end{cases}$ (15)
where $t\in[T]$ and $\tau$ is uniformly chosen from $\{1,2,\dots,\lfloor\beta T\rfloor\}$ with $\beta=\frac{2}{3}$. These three patterns have variations bounded by $1$.
To generate opponents’ highest bid sequences with larger variations, we partition the time horizon $T$ into $\lceil V_{T}\rceil$ segments, each of length at least $3$, and apply one of the three building blocks from Equation (15) to each segment. The fourth pattern is generated by the sinusoidal wave $m_{t}=\frac{1}{2}+\frac{1}{2}\sin\left(\frac{V_{T}\pi t}{T}\right)$, which is employed in Besbes et al. (2019), Cheung et al. (2022).
Experimental setup: We choose $T\in\{5000,8000,\dots,59000\}$ and let $V_{T}=\frac{1}{4}\cdot T^{\alpha}$ for $\alpha\in\{0.1,0.3,0.5,0.7,0.9\}$. The learner’s values $(v_{t})_{t=1}^{T}$ are drawn i.i.d. from the uniform distribution on $[0,1]$. We consider the following algorithms:
  • Theoretical Bound: The theoretical dynamic regret upper bound $O\left(\sqrt{TV_{T}}\ln T\right)$.
  • AR-Prod: Algorithm 2 for unknown $V_{T}$. We use $\eta=1$, $\epsilon=\frac{4}{\sqrt{T}}$, and $c=\frac{1}{T}$.
  • AR-OMD: Algorithm 3 for unknown $L_{T}$. We use $\epsilon=\frac{1}{T^{0.9}}$, $\eta=\sqrt{\ln(T^{0.9})}$, and we consider $m_{t}\neq m_{t-1}$ if $|m_{t}-m_{t-1}|\geq 10^{-6}$.
  • BOBW: Algorithm 4 with best-of-both-worlds guarantee. We use the default parameters specified in Algorithm 4.
Figure 2: Slopes of dynamic regret rates against varying time horizons for different opponents’ highest bid sequence patterns and values of $\alpha$.
Results: For each value of $\alpha$, we compute the slopes of log-log plots of average dynamic regret against $T$, and present all results in Figure 2. For all four temporal patterns, we observe that the slopes of AR-Prod align closely with the theoretical upper bound, which is consistent with Theorem 4.6. For the “Constant” pattern, AR-OMD’s slope approaches $\alpha$ because this pattern implies $V_{T}=L_{T}$ and its theoretical dynamic regret bound is $O\left(L_{T}\sqrt{\ln T}\right)=O\left(T^{\alpha}\sqrt{\ln T}\right)$ by Theorem 4.9. For the remaining three patterns, the slopes of AR-OMD are close to $1$. This is because for these three continuously evolving patterns, $L_{T}=\Omega(T)$ even when $V_{T}$ is very small. Finally, BOBW’s slope is consistently the minimum of the AR-Prod and AR-OMD slopes, confirming Theorem 4.10.
6.2 Performance Against Budget-Pacing Bidders
We evaluate our algorithms against opponents using the budget-pacing algorithm from Gaitonde et al. (2022). We consider this experimental setting for two reasons: (i) the budget-pacing policy is an important strategy in both second-price auctions (Balseiro and Gur 2019) and first-price auctions (Gaitonde et al. 2022); (ii) when the value of each opponent varies slowly, the sequence of opponents’ highest bids is also slowly varying. We find that our algorithms outperform the baselines, especially in regimes where opponents have limited budgets, even when the sequence of opponents’ values varies rapidly.
For the budget-pacing algorithm from Gaitonde et al. (2022), the pacing multiplier $\mu_{t}$ is adaptively adjusted based on the observed expenditure to maintain a target spending rate $\rho_{k}=B_{k}/T$, where $k$ is the index of the $k$-th agent and $B_{k}$ is the initial budget of the $k$-th agent. We set $\epsilon_{k}=\frac{1}{\sqrt{T}}$ and $\bar{\mu}=\frac{T}{B_{k}}-1$ for this experiment based on suggestions from Gaitonde et al. (2022).
Slowly Varying Property: When opponent bidders have slowly varying valuations, the resulting sequence of opponents’ highest bids is also slowly varying. This occurs because, given the current budget, Lagrangian multiplier, and ad impression value, each budget-pacing bidder’s bid is precisely determined. This makes budget-pacing a suitable opponent strategy for validating our algorithms’ performance in practical scenarios.
Experimental setup: We consider two budget regimes with 20 opponents:
  • Sufficient budget: Each opponent has budget $T/20$
  • Insufficient budget: Each opponent has budget $T/40$
The sufficient budget regime corresponds to the case where the combined budget of all opponents is sufficient to purchase every ad impression, potentially leaving nothing for the learner. In contrast, the insufficient budget regime refers to the case where the opponents’ budgets are collectively insufficient to purchase every ad impression.
The learner’s values are i.i.d. uniform on $[0,1]$, while opponent values follow the four patterns from Section 6.1, scaled by $0.8$ to ensure a reasonable winning probability for the learner. We set $T=12000$, $V_{T}=\frac{1}{4}\cdot T^{\alpha}$ with $\alpha\in\{0,0.1,\dots,1\}$, and average results over 50 runs. Baselines include the Hedge algorithm and the SEW algorithm (Han et al. 2020)  [^2] We use the official implementation of the SEW policy (Han 2024) in our numerical simulations..
Results: Figures 3 and 4 show that our algorithms, particularly AR-Prod, outperform baselines in both budget regimes, even in the case of $\alpha=1$. The advantage is more pronounced in the insufficient budget regime, where opponents bid more conservatively due to lower target spend rates, creating slowly varying opponents’ highest bid sequences that our algorithms can exploit.
Figure 3: Cumulative rewards against budget-pacing bidders in the sufficient budget regime.
Figure 4: Cumulative rewards against budget-pacing bidders in the insufficient budget regime.
Robustness analysis: To ensure that our advantage arises from algorithmic adaptivity rather than simply from exploiting slowly varying sequences, we conducted additional experiments where both the learner’s and the opponents’ values are drawn from the same distributions: uniform, a truncated Gaussian (with mean $0.4$ and standard deviation $0.2$), or Beta(3,3)(3,3). In these experiments, we set $T=12000$ with $20$ budget-pacing bidders and vary the initial budget of each bidder. We repeat each experiment $50$ times and report the average performance. As shown in Figure 5, our algorithms remain competitive in regimes with sufficient budgets and outperform baselines in regimes with insufficient budgets, confirming that their adaptive capabilities extend beyond merely capitalizing on slow variation.
Figure 5: Cumulative rewards against budget-pacing bidders for different initial budgets.
7 Conclusion
This work examines online first-price auctions within non-stationary environments. While prior research typically focuses on competing against the best fixed policy in hindsight, such a policy can be suboptimal even in environments with mild non-stationarity. We instead investigate conditions under which competition against a dynamic benchmark, achieving the highest possible revenue, is feasible. We identify two measures of regularity on the opponents’ highest bid sequence and establish minimax-optimal dynamic regret rates for the class of auction instances where the sequence of opponents’ highest bids satisfies either of these regularity constraints. For future work, it would be valuable to investigate tight dynamic regret rates under settings where only winning-bid feedback or binary feedback is available. From a technical perspective, our analysis considers the dynamic regret of a specific one-sided Lipschitz function with a single discontinuity. Given the existence of important one-sided Lipschitz functions with multiple discontinuities (Dütting et al. 2023), investigating the applicability of our algorithms to these more general settings presents a compelling research direction.
[…]
We thank Yanjun Han from New York University for providing the code of the SEW policy. This research is generously supported by the NSF grant CCF-2106508.